OpenStudy (anonymous):

The regular retail price of a particular box of a well-known brand of cereal ranges from $2.80 to $3.25 across the grocery stores in the city. The distribution of the prices follows a uniform distribution. If 12 stores are selected at random and the regular retail price of the cereal is recorded, what is the probability exactly 5 of the stores charge less than $3.00?

OpenStudy (anonymous):

Is this a poisson distribution?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i'm not sure on this. @callisto

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if it uniform all prices are equally likely. so the prob of one store charging less than 3.00 is 20 out 46

OpenStudy (anonymous):

no it is 0.22

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

i did not get it. That is the answer.

OpenStudy (anonymous):

so now it a bionomial which is very similar to the poission dist. 12C5*(20/46)^5*(26/46)^7

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%2820%2F46%29%5E5*%2826%2F46%29%5E7*12choose5 according to wolfram is .22267

OpenStudy (anonymous):

okay. Thanks.

OpenStudy (anonymous):

can you do it using the a possion dist?

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

we have to decide what alpha is first

OpenStudy (anonymous):

hmm... i see.

OpenStudy (anonymous):

the average for the whole population is the same as the unknown small sample ie 12

OpenStudy (anonymous):

i know that part. but how should we figure out?

OpenStudy (anonymous):

ahh for example if i said half the shops were less than average.. 12*.5=6 then alpha would be 6 but were dealing with 3.00 so its 20/46 12*(20/46)=5.2173913 so ALPHA is 5.2173913

OpenStudy (anonymous):

how did you get 20/46?

OpenStudy (anonymous):

pick a number between 0 and 5 ... there is 6 choices even though 5-0=5 so with 3.25-2.80=45cents there is 46 choices now less than $3 is 2.99 to 2.80 2.99-2.80=20 choices so prob(x<$3)=20/46

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

what uniform is saying is every price has the same prob of occuring P(x=2.80)=P(x=2.81)=....................................=P(x=3.25) so 46 equal likely probs which all add to 1 1/46 is there prob individually

OpenStudy (anonymous):

so we worked out alpha = 5.2173913 and x=5

OpenStudy (anonymous):

sub into the possion equation and post your answer

OpenStudy (anonymous):

0.6160?

OpenStudy (anonymous):

this question is tough.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=5.2173913%5E5*exp%28-5.2173913%29%2F5%21 this gives the answer as 0.1746633

OpenStudy (anonymous):

which like we said earlier the possion dist is like bionomial.... in that its an estimate.. the possion dist works better for large numbers

OpenStudy (anonymous):

goodnight

OpenStudy (anonymous):

ok. good night